The AVC of producing an instrument comes to Rs.2P units and Fixed Cost of Rs.90000. Find the Break even point if the instrument sells for Rs.20 per unit.
1
Expert's answer
2015-09-02T10:55:10-0400
The break-even level or break-even point (BEP) represents the sales amount—in either unit or revenue terms—that is required to cover total costs (both fixed and variable). Total profit at the break-even point is zero.
X = TFC/(P - AVC) = 90000/(20 - 0.2*20) = 90000/16 = Rs 5625